OpenStudy (anonymous):

f(x)=-2x-7 and g(x)=-4x+6. Find the (g o f)(-5)

hartnn (hartnn):

you know how to get (g o f) (x) ??

OpenStudy (anonymous):

i do not remember

hartnn (hartnn):

ok, to get (g o f) (x) just plug in 'x' = -2x-7 (which is f(x)) in place of 'x' in -4x+6 (which is g(x))

OpenStudy (anonymous):

so -4x+6(-2x-7))?

hartnn (hartnn):

you just multiplied them ? you need to replace the x in g(x) g o f (x) = -4 (-2x-7) +6 = ... ?

OpenStudy (anonymous):

do i distribute the -4 into the parentheses

hartnn (hartnn):

yes! do it...

OpenStudy (anonymous):

8x+28=6

hartnn (hartnn):

you mean 8x+28+6 ?

OpenStudy (anonymous):

yeah

hartnn (hartnn):

then, yes that would be your (g o f)(x) to get (g o f) (-5), just plug in x=-5 in that!

OpenStudy (anonymous):

8(-5)+28+6=-6

hartnn (hartnn):

correct!

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^